A particle has such a curvilinear motion that its x coordinate is defined by π‘₯ = 5𝑑 3 βˆ’105𝑑 where π‘₯ is in inches and 𝑑 is in seconds. When 𝑑 = 2 𝑠, the total acceleration is 75 𝑖𝑛/𝑠 2 . If the y component of acceleration is constant and the particle starts from rest at the origin when 𝑑 = 0, determine its total velocity when 𝑑 = 4 𝑠.


Answer: 𝒗 = πŸπŸ–. πŸ•πŸ“ 𝒇𝒕/s


1
Expert's answer
2021-02-25T04:15:15-0500

Total acceleration , "\vec{a}=a_x{i}+a_y{j}=30t{i}+a_y{j}"

"\vec{a}= \sqrt{900t^2+a_y^2}; t=2 \implies75=\sqrt{900*2^2+a_y^2} \implies a_y=45 in/s^2"

"\vec{a}=[30t{i}+45{j}] in/s^2"

Velocity at t=4 sec "\implies \vec{v}=\vec{u}+\vec{a}t=-105i+ [(30*4){i}+45{j}]*4"

"\vec{v}=[375{i}+180{j}]in/s"

"\vec{|v|}=\sqrt{375^2+180^2}=415.96 in/s=415.96in/s^2=34.66 ft/s^2"